Substance 18. Amorphophallus johnsonii is a plant
Please show every step and calculation

Substance 18. Amorphophallus johnsonii is a plant
Substance 18. Amorphophallus johnsonii is a plant
Please show every step and calculation
Answer #1

A) MEAN

Step 1: write all the values (X)

Value (X)

51

45

61

76

11

117

7

132

52

149

Step 2: Calculate the sum of numbers (SX)

SX = 701

Step 3: calculate Mean using the following formula,
where, n = total number of X

Mean = SX/n

Mean = 701/10

Mean = 70.1

A) STANDARD DEVIATION

Step 1

Step 2

Step 3

X

X-Mean

Square of X-Mean

51

-19.1

364.81

364.81

45

-25.1

630.01

630.01

61

-9.1

82.81

82.81

76

5.9

34.81

34.81

11

-59.1

3492.81

3492.81

117

46.9

2199.61

2199.61

7

-63.1

3981.61

3981.61

132

61.9

3831.61

3831.61

52

-18.1

327.61

327.61

149

78.9

6225.21

6225.21

21170.9

Step 4. Sum of Square of X-Mean

S(X-Mean) = 21170.9

Step 5: Calculate variance using following formula

Variance = S(X-Mean)/N

Variance = 21170.9/10

Variance = 2117.09

Step 6. Standard Deviation is square root of variance

46.01185

B) STANDARD ERROR

Calculate standard error using following formula:

standard error = Standard deviation/sqrt n

standard error = 46.01185/sqrt 10

standard error = 46.01185/3.16

standard error = 14.55

C)

CI = Mean ± Standard error

CI = 70.1 - 14.55

CI = 55.55

CI = 70.1 + 14.55

CI = 84.65

D) If we increase the number set from 10 to 25, then that will change the mean of the entire set, which will change the deviation from the mean for each number on the list, which changes the standard deviation. If you include an additional number or a few additional numbers that are far away from the other numbers, this inclusion will increase the standard deviation.